# Euclid: Early Release Observations – The star cluster systems of the Local Group dwarf galaxies IC 10 and NGC 6822††thanks: This paper is published on behalf of the Euclid Consortium

### v1

@@ 7.3 Luminosity and colour functions
The left panel of Fig. 14 shows the star cluster luminosity functions (LFs), expressed in extinction-corrected absolute V-band magnitude calculated using the BAGPIPES outputs for and the distances in Table 1. We show the overall LFs for each galaxy, as well as the LFs for the young and old clusters using the definitions introduced in Sect. 7.2. For completeness, we also add the NGC 6822 GCs SC1, SC2, and SC4, which lie outside the Euclid FoV, using the magnitudes and colours from Veljanoski et al. (2015).
@@
The right-hand side of Fig. 14 shows the colour distributions. Our old clusters are found within the colour range spanned by the Georgiev et al. (2009) GC sample, while most of our young clusters are found blueward of our old clusters. The colours of the old clusters peak at a similar value compared to the GCs in the Georgiev et al. (2009) sample in both galaxies, with a peak around . Georgiev et al. (2009) highlight an absence of faint () blue GCs in their sample, defined as those with . We have several old clusters that fit these criteria in both galaxies (with ages between 3 and 12 Gyr), but we note that the uncertain reddening may be impacting these colours.
@@ 7.4 Half-light radii
It is interesting to explore the correlation between the half-light radii of the old clusters and their luminosities. Since changes very little over many relaxation times (e.g., Meylan & Heggie 1997), this can potentially provide insight into the formation process of the most massive, long-lived clusters. Figure 15 shows how the half-light radii of the old clusters depend on absolute V-band magnitude. In the MW, GCs show a clear trend in with , in the sense that more luminous clusters are more compact. Very few clusters exist above the empirical boundary (Mackey & van den Bergh 2005), which is reproduced in Fig. 15. For comparison, the Georgiev et al. (2009) sample is also plotted. Our old clusters lie below the empirical line and fall within the size range of the Georgiev et al. (2009) sample. Their half-light radii also peak at a similar value to ours (around pc), as well as the old [] clusters in the Gatto et al. (2021) sample. The brighter old clusters () have half-light radii within the range –4.5 pc, similar to what is seen in MW GCs (Harris 1996). However, toward fainter magnitudes, the spread increases considerably for NGC 6822, reaching up to 12 pc, and also extending beyond the half-light radii of the Georgiev et al. (2009) sample at these faint magnitudes. This increased scatter at faint magnitudes has also recently been observed in NGC 2403 and IC 342 using Euclid data (Larsen et al. 2025). On the other hand, the size distribution of old clusters in IC 10 is relatively flat as a function of magnitude. In both NGC 6822 and IC 10, there is a subtle trend where clusters located farther from the galaxy centre tend to have slightly larger half-light radii.
@@
Figure 15 shows that NGC 6822 has three particularly extended old clusters (red circles), with pc, which have no counterparts in IC 10. While a few examples old extended clusters exist in the Georgiev et al. (2009) sample, they are all brighter than the ones highlighted here. Two of the extended clusters in NGC 6822, SC3 and SC5, were previously known from ground-based work (Hwang et al. 2011; Huxor et al. 2013) while the other is newly identified in this work. With pc, ESCC-NGC6822-27 is the most extended old cluster within our sample, and it has an age of Gyr, a mass of and a metallicity of . Thumbnails of the extended clusters within our FoV are shown in Fig. 16, where their very diffuse nature is readily apparent. Another three extended clusters are known in NGC 6822 which lie beyond the Euclid FoV (Hwang et al. 2011; Huxor et al. 2013). A possible explanation for the origin of extended clusters is that they have been born extended and evolved in a weak tidal field (Hurley & Mackey 2010; Bianchini et al. 2014), consistent with their existence in the outskirts of a dwarf galaxy such as NGC 6822. The fact that extended clusters are also seen in the halos of massive galaxies (e.g., Huxor et al. 2005; Mouhcine et al. 2010; Jang et al. 2012) provides tantalising evidence that they have been donated via dwarf galaxy accretion events.
@@
We have also searched for correlations between mass and size, and age and size within our sample. A modest but statistically significant correlation between cluster mass and size exists in NGC 6822, but not in IC 10. In contrast, no clear relationship is found between cluster age and size in either galaxy. While some studies have seen evidence for such trends (e.g. Lee et al. 2005; Chandar et al. 2016; Mackey & Gilmore 2003; Ryon et al. 2015; Bastian et al. 2012; Scheepmaker et al. 2007), others do not (e.g. Larsen 2004; Barmby et al. 2009). Overall, our findings suggest that while mass may influence cluster size to some extent, age does not appear to be a strong driver.
@@ 7.5 Cluster metallicities, extinctions, and ages
Figure 17 shows how the metallicity of the clusters varies as a function of age, using the BAGPIPES outputs. In addition to the individual cluster values (light grey), we also show the mean metallicity in specific age bins (black points), along with its standard deviation. The age bins correspond to our previous definitions of young, intermediate, and old populations, with the intermediate-age category further divided into two bins. Both galaxies show the expected behaviour, with the most recently formed clusters having the highest metallicities and older clusters having lower metallicities. Beyond the first age bin, the age-metallicity relation is flat to within the uncertainties in both systems, with metallicities at a given age being slightly higher in IC 10 than in NGC 6822. Within any age bin, there is significant scatter when examining individual clusters, only some of which is likely to be real. Indeed, the plots also hint at the existence of potential systematics: a preference for low-metallicity fits reaching the floor of the prior in the second age bin, as well as some unexpectedly high-metallicity fits in the oldest age bins. In the latter case, we checked that these clusters did not show a correlation with far-IR dust emission (by inspecting SPIRE 500 and Spitzer 160 images), or with strong sources; this confirmed that their SEDs were not contaminated by non-stellar emission. One possible explanation for the high metallicities could be due to our extinction priors. These were set by considering the range of reddening values in the published literature; however, these may not capture the full range of reddening across the Euclid FoV. The Schlegel et al. (1998) dust maps, recalibrated by Schlafly & Finkbeiner (2011), suggest higher values in some places than our adopted limits. If the true extinction is higher than our allowed range, the SED fitting may only be able to reproduce the red colours by favouring older ages and higher metallicities. Therefore, to evaluate the consistency of our extinction estimates, we converted our to , assuming , and compared them to the high-resolution reddening map derived from resolved RGB stars in Hunt et al. (2025, Annibali et al., in prep.), who inferred by matching each stars location in the uncorrected vs. CMD to a grid of reference RGB lines with varying reddening. We find that the mean offset is minimal at , with a standard deviation of 0.14, suggesting that our adopted extinction prior is appropriate despite the spatial variation across the field. Additionally, when we investigated removing the upper limit on the extinction prior, the clusters still broadly occupied the same region. Alternatively, stochastic effects could be responsible for assigning old ages and high metallicities to these clusters (Appendix A). While Fig.17 illustrates the qualitative trends in the metallicity evolution of the galaxies, a more detailed reconstruction of the age-metallicity relationship of the star clusters would require either CMD fitting or spectroscopy.
@@ 7.6 Globular cluster scaling relations
The GC populations of dwarf galaxies are of particular interest given their likely role in building up the halos of more massive systems. Furthermore, a remarkable diversity in the GC populations of low-mass galaxies has been uncovered in recent years (e.g., Eadie et al. 2022; Jones et al. 2023; Saifollahi et al. 2025) and it is interesting to explore how NGC 6822 and IC 10 compare.
Unfortunately, there is no universally accepted definition of what constitutes a GC (see Gratton et al. 2019). In the MW and its satellites, GCs are generally identified as ancient halo, bulge and disc star clusters. On the other hand, accurate age information is rarely available in extragalactic studies, where we also only have knowledge of projected positions. Instead, extragalactic GC candidates are generally identified on the basis of magnitude, colour, and size information, using stellar population models to select clusters older than a few gigayears (e.g., Carlsten et al. 2022; Floyd et al. 2024; Lim et al. 2025). To define a sample of GCs that bridges both the MW and extragalactic conventions, we use the results of our visual classification along with the output of BAGPIPES. Specifically, we adopt the following criteria:
@@
In dwarf galaxies, GCs are typically metal poor (e.g., Larsen et al. 2022), consistent with the mass-metallicity relation. For galaxies with stellar masses like NGC 6822 and IC 10, simulations suggest that star clusters formed 5 Gyr ago would be expected to have (Kruijssen 2019; Ma et al. 2016) and so we apply a generous cut to exclude more metal-rich clusters. There is no well-established age criterion for a GC. In the MW, most GCs are Gyr old, but there are also halo star clusters that are several gigayears younger (e.g., Weisz et al. 2016). Along with the metallicity cut, the age cut of 5 Gyr is designed to capture all long-lived metal-poor star clusters without being overly restrictive. We note that our selection criteria would retain all but one of the six GCs in the Fornax dSph galaxy. The excluded object would be Fornax 6, which has a spectroscopically derived metallicity of and an estimated age of Gyr from Gaia photometry and MIST isochrone fitting (Pace et al. 2021). Given its young age and relatively high metallicity, Fornax 6 is arguably not what is typically considered to be a GC.
@@
Figure 11 shows that the age cut adopted guarantees that all clusters are more massive than a few times . With these criteria, we arrive at 7 GCs in NGC 6822, and 8 GCs in IC 10. Of the star clusters previously identified as NGC 6822 GCs, Hubble-VII, SC6, and 7 fall within our definition but SC3 does not. The SED-fitting routine assigns a rather young age (2.5 Gyr) and relatively high metallicity () to this cluster, which is the faintest of the known GCs in NGC 6822. Based on a previous spectroscopic analysis of this object (Hwang et al. 2014, see Table 7), as well as analysis of a deep HST CMD (McGill et al. 2025), we believe the BAGPIPES fit is unreliable and that the cluster is older and more metal poor; we therefore include SC3 in the total number of NGC 6822 GCs. We further add in SC5, for which we could not perform SED-fitting due to lack of coverage, as well as SC1, SC2, and SC4, all of which lie outside the Euclid FOV. This brings the total number of GCs in NGC 6822 to 11. The resulting GC candidates are flagged in the thumbnails shown in Appendix C and the spatial distribution of those that fall within the Euclid FoV are displayed as black points in Fig. 13. In both galaxies, the GCs exhibit extended distributions – in NGC 6822, 7 of the 11 GCs lie outside the galaxy half-light radius (indicated by dashed grey ellipse) while all 8 of the GCs in IC 10 lie outside of it. Interestingly, NGC 6822’s GCs appear to be preferentially located on the eastern side of the galaxy (this remains true even when considering the GCs outwith the Euclid footprint). The most obvious object for comparison is the SMC which, due to its proximity, has many cluster age and metallicity determinations from CMD fitting (e.g. Glatt et al. 2008). The SMC is generally considered to have only one old GC, NGC 121 (e.g., Forbes et al. 2018b)999We note that two faint GC candidates have recently been discovered at large radius in the SMC but their true nature is not yet clear (Cerny et al. 2021, 2023).. We use the Bica et al. (2020) catalogue to search for high-confidence star clusters (object class “C”) that satisfy our metallicity and age criteria. We find 14 SMC clusters fall within our definition of GC, which is slightly larger than the numbers found for NGC 6822 and IC 10.
@@
The specific frequency of GCs, defined as the number of GCs per unit of galaxy luminosity, was introduced to measure the richness of GC systems in galaxies (Harris & van den Bergh 1981) and remains commonly used today. It is calculated as , where is the absolute V-band magnitude of the host galaxy and is the number of GCs. The values are impacted by the galaxy’s recent star formation history, affecting its , and the dynamical evolution history of its GCs, affecting its . Using the galaxy absolute magnitudes from Table 1, and of 11 and 8, we find that is and in NGC 6822 and IC 10, respectively. The uncertainty in reflects both Poisson noise in the GC counts and the reported uncertainty in . These values are displayed in the top panel of Fig. 19, plotted against . This plot also shows and values for the sample of LG and isolated late-type dwarf galaxies studied by Forbes et al. (2018b), which includes some of the systems originally presented in Georgiev et al. (2010). We add further datapoints from the recent studies by Karim et al. 2024 (IC 2574) and Larsen et al. 2025 (Holmberg II), and we include values for the SMC calculated assuming both one and 14 GCs (connected via a dashed line). We see that NGC 6822 and IC 10 have high values for their luminosity, and lie at the high end of the distribution when compared to the sample of dwarfs as a whole. This is particularly true when considering that this plot only shows dwarf galaxies that have at least one identified GC; many others show none at all. Of the three late-type galaxies with higher values, one of them is the LG transition dwarf galaxy Pegasus (Cole et al. 2017) while the other two are more distant systems (Georgiev et al. 2010, KK 16 and UGC 685).
@@ 8 Conclusions
The clusters identified have masses spanning from to , and span ages from 10 Myr to 10 Gyr. IC 10 has more young clusters than NGC 6822, and its young clusters also reach higher masses (), consistent with this galaxy having a higher truncation mass on account of its higher SFR surface density. We found several examples of old massive () clusters in both dwarfs that could be the descendants of the young star clusters recently uncovered in low-mass galaxies at high redshift (e.g., Mowla et al. 2024; Adamo et al. 2024). We highlighted the existence of a particularly exceptional object in the outskirts of NGC 6822, SC7, for which we made the first mass determination of , rivalling the most massive GC in the MW. Combined with its high ellipticity and moderately high metallicity, this object bears many hallmarks of being the accreted core of a now-defunct smaller system.
@@
Lastly, we used well-defined criteria to select a subsample of clusters as candidate GCs; this yields 11 objects in NGC 6822 (folding in objects from previous studies) and eight in IC 10. The GCs are more broadly distributed than other star clusters, with 7 of the 11 GCs lying outside the galaxy half-light radius in NGC 6822 and all 8 of the GCs in IC 10 lying outside of it. We placed the systems on the well-known GC scaling relations and found that both galaxies have high values for their luminosity, and lie at the high end of the distribution when compared to the population of dwarfs as a whole. We note that although the new star clusters uncovered in this work have increased the GC count in both galaxies, more GCs may still await detection. Indeed, the Euclid ERO images do not cover the entire halos of these systems, where more GCs may reside (including 3 previously known GCs in NGC 6822). This highlights the need for systematic, high-sensitivity, wide-field surveys of dwarf galaxies to obtain an accurate census of their GC populations.

### v3

@@ 7.3 Luminosity and colour functions
The left panels of Fig. 14 show the star cluster luminosity functions (LFs), expressed in extinction-corrected absolute V-band magnitude calculated using the BAGPIPES outputs for and the distances in Table 1. We show the overall LFs for each galaxy, as well as the LFs for the young and old clusters using the definitions introduced in Sect. 7.2. For completeness, we also add the NGC 6822 GCs SC1, SC2, and SC4, which lie outside the Euclid FoV, using the magnitudes and colours from Veljanoski et al. (2015).
@@
The right-hand panels of Fig. 14 show the colour distributions. Our old clusters are found within the colour range spanned by the Georgiev et al. (2009) GC sample, while most of our young clusters are found blueward of our old clusters. The colours of the old clusters peak at a similar value compared to the GCs in the Georgiev et al. (2009) sample in both galaxies, with a peak around . Georgiev et al. (2009) highlight an absence of faint () blue GCs in their sample, defined as those with . We have several old clusters that fit these criteria in both galaxies (with ages between 3 and 12 Gyr), but we note that the uncertain reddening may be impacting these colours.
@@ 7.4 Half-light radii
It is interesting to explore the correlation between the half-light radii of the old clusters and their luminosities. Since changes very little over many relaxation times (e.g. Meylan and Heggie 1997), this can potentially provide insight into the formation process of the most massive, long-lived clusters. Figure 15 shows how the half-light radii of the old clusters depend on absolute V-band magnitude. In the MW, GCs show a clear trend in with , in the sense that more luminous clusters are more compact. Very few clusters exist above the empirical boundary (Mackey and van den Bergh 2005), which is reproduced in Fig. 15. For comparison, the Georgiev et al. (2009) sample is also plotted. Our old clusters lie below the empirical line and fall within the size range of the Georgiev et al. (2009) sample. Their half-light radii also peak at a similar value to ours (around pc), as well as the old [] clusters in the Gatto et al. (2021) sample. The brighter old clusters () have half-light radii within the range –4.5 pc, similar to what is seen in MW GCs (Harris 1996). However, toward fainter magnitudes, the spread increases considerably for NGC 6822, reaching up to 12 pc, and also extending beyond the half-light radii of the Georgiev et al. (2009) sample at these faint magnitudes. This increased scatter at faint magnitudes has also recently been observed in NGC 2403 and IC 342 using Euclid data (Larsen et al. 2025). On the other hand, the size distribution of old clusters in IC 10 is relatively flat as a function of magnitude. In both NGC 6822 and IC 10, there is a subtle trend where clusters located farther from the galaxy centre tend to have slightly larger half-light radii.
@@
Figure 15 shows that NGC 6822 has three particularly extended old clusters (red circles), with pc, which have no counterparts in IC 10. While a few examples old extended clusters exist in the Georgiev et al. (2009) sample, they are all brighter than the ones highlighted here. Two of the extended clusters in NGC 6822, SC3 and SC5, were previously known from ground-based work (Hwang et al. 2011; Huxor et al. 2013) while the other is newly identified in this work. With pc, ESCC-NGC6822-27 is the most extended old cluster within our sample, and it has an age of Gyr, a mass of and a metallicity of . Thumbnails of the extended clusters within our FoV are shown in Fig. 16, where their very diffuse nature is readily apparent. Another three extended clusters are known in NGC 6822 which lie beyond the Euclid FoV (Hwang et al. 2011; Huxor et al. 2013). A possible explanation for the origin of extended clusters is that they have been born extended and evolved in a weak tidal field (Hurley and Mackey 2010; Bianchini et al. 2014), consistent with their existence in the outskirts of a dwarf galaxy such as NGC 6822. The fact that extended clusters are also seen in the halos of massive galaxies (e.g. Huxor et al. 2005; Mouhcine et al. 2010; Jang et al. 2012) provides tantalising evidence that they have been donated via dwarf galaxy accretion events.
@@
We have also searched for correlations between mass and size, and age and size within our sample. A modest but statistically significant correlation between cluster mass and size exists in NGC 6822, but not in IC 10. In contrast, no clear relationship is found between cluster age and size in either galaxy. While some studies have seen evidence for such trends (e.g. Lee et al. 2005; Chandar et al. 2016; Mackey and Gilmore 2003; Ryon et al. 2015; Bastian et al. 2012; Scheepmaker et al. 2007), others do not (e.g. Larsen 2004; Barmby et al. 2009). Overall, our findings suggest that while mass may influence cluster size to some extent, age does not appear to be a strong driver.
@@ 7.5 Cluster metallicities, extinctions, and ages
Figure 17 shows how the metallicity of the clusters varies as a function of age, using the BAGPIPES outputs. In addition to the individual cluster values (light grey), we also show the mean metallicity in specific age bins (black points), along with its standard deviation. The age bins correspond to our previous definitions of young, intermediate, and old populations, with the intermediate-age category further divided into two bins. Both galaxies show the expected behaviour, with the most recently formed clusters having the highest metallicities and older clusters having lower metallicities. Beyond the first age bin, the age-metallicity relation is flat to within the uncertainties in both systems, with metallicities at a given age being slightly higher in IC 10 than in NGC 6822. Within any age bin, there is significant scatter when examining individual clusters, only some of which is likely to be real. Indeed, the plots also hint at the existence of potential systematics: a preference for low-metallicity fits reaching the floor of the prior in the second age bin, as well as some unexpectedly high-metallicity fits in the oldest age bins. In the latter case, we checked that these clusters did not show a correlation with far-IR dust emission (by inspecting SPIRE 500 and Spitzer 160 images), or with strong sources; this confirmed that their SEDs were not contaminated by non-stellar emission. One possible explanation for the high metallicities could be due to our extinction priors. These were set by considering the range of reddening values in the published literature; however, these may not capture the full range of reddening across the Euclid FoV. The Schlegel et al. (1998) dust maps, recalibrated by Schlafly and Finkbeiner (2011), suggest higher values in some places than our adopted limits. If the true extinction is higher than our allowed range, the SED fitting may only be able to reproduce the red colours by favouring older ages and higher metallicities. We explored this by assessing the impact of changing the dust prior in the BAGPIPES fitting process. First, we removed the upper limit on the dust prior and found the old, high-metallicity clusters still broadly occupied the same region. We then replaced the global dust prior for the entire galaxy with a local one specific to each cluster. For this, we exploited the high-resolution reddening map derived from resolved RGB stars (Annibali et al., in prep), who inferred by comparing the CMD positions of stars to a fiducial RGB sequence. Setting the prior for each cluster as a gaussian of width 0.2 centered on the local , we redid the fits but found very little change to the results, with the old high-metallicity fits persisting. Alternatively, stochastic effects could be responsible for assigning old ages and high metallicities to these clusters (Appendix A). While Fig.17 illustrates the qualitative trends in the metallicity evolution of the galaxies, a more detailed reconstruction of the age-metallicity relationship of the star clusters would require either CMD fitting or spectroscopy.
@@ 7.6 Globular cluster scaling relations
The GC populations of dwarf galaxies are of particular interest given their likely role in building up the halos of more massive systems. Furthermore, a remarkable diversity in the GC populations of low-mass galaxies has been uncovered in recent years (e.g. Eadie et al. 2022; Jones et al. 2023; Saifollahi et al. 2025) and it is interesting to explore how NGC 6822 and IC 10 compare.
Unfortunately, there is no universally accepted definition of what constitutes a GC (see Gratton et al. 2019). In the MW and its satellites, GCs are generally identified as ancient halo, bulge and disc star clusters. On the other hand, accurate age information is rarely available in extragalactic studies, where we also only have knowledge of projected positions. Instead, extragalactic GC candidates are generally identified on the basis of magnitude, colour, and size information, using stellar population models to select clusters older than a few gigayears (e.g. Carlsten et al. 2022; Floyd et al. 2024; Lim et al. 2025). To define a sample of GCs that bridges both the MW and extragalactic conventions, we use the results of our visual classification along with the output of BAGPIPES. Specifically, we adopt the following criteria:
@@
In dwarf galaxies, GCs are typically metal poor (e.g. Larsen et al. 2022), consistent with the mass-metallicity relation. For galaxies with stellar masses like NGC 6822 and IC 10, simulations suggest that star clusters formed 5 Gyr ago would be expected to have (Kruijssen 2019; Ma et al. 2016) and so we apply a generous cut to exclude more metal-rich clusters. There is no well-established age criterion for a GC. In the MW, most GCs are Gyr old, but there are also halo star clusters that are several gigayears younger (e.g. Weisz et al. 2016). Along with the metallicity cut, the age cut of 5 Gyr is designed to capture all long-lived metal-poor star clusters without being overly restrictive. We note that our selection criteria would retain all but one of the six GCs in the Fornax dSph galaxy. The excluded object would be Fornax 6, which has a spectroscopically derived metallicity of and an estimated age of Gyr from Gaia photometry and MIST isochrone fitting (Pace et al. 2021). Given its young age and relatively high metallicity, Fornax 6 is arguably not what is typically considered to be a GC.
@@
Figure 11 shows that the age cut adopted guarantees that all clusters are more massive than a few times . With these criteria, we arrive at 6 GCs in NGC 6822, and 9 GCs in IC 10. Of the star clusters previously identified as NGC 6822 GCs, Hubble-VII, SC6, and 7 fall within our definition but SC3 does not. The SED-fitting routine assigns a rather young age (2.5 Gyr) and relatively high metallicity () to this cluster, which is the faintest of the known GCs in NGC 6822. Based on a previous spectroscopic analysis of this object (Hwang et al. 2014, see Table 7), as well as analysis of a deep HST CMD (McGill et al., in prep), we believe the BAGPIPES fit is unreliable and that the cluster is older and more metal poor; we therefore include SC3 in the total number of NGC 6822 GCs. We further add in SC5, for which we could not perform SED-fitting due to lack of coverage, as well as SC1, SC2, and SC4, all of which lie outside the Euclid FOV. This brings the total number of GCs in NGC 6822 to 11. The resulting GC candidates are flagged in the thumbnails shown in Appendix C and the spatial distribution of those that fall within the Euclid FoV are displayed as black points in Fig. 13. In both galaxies, the GCs exhibit extended distributions – in NGC 6822, 7 of the 11 GCs lie outside the galaxy half-light radius (indicated by dashed grey ellipse) while all 9 of the GCs in IC 10 lie outside of it. Interestingly, NGC 6822’s GCs appear to be preferentially located on the eastern side of the galaxy (this remains true even when considering the GCs outwith the Euclid footprint). The most obvious object for comparison is the SMC which, due to its proximity, has many cluster age and metallicity determinations from CMD fitting (e.g. Glatt et al. 2008). The SMC is generally considered to have only one old GC, NGC 121 (e.g. Forbes et al. 2018b)999We note that two faint GC candidates have recently been discovered at large radius in the SMC but their true nature is not yet clear (Cerny et al. 2021, 2023).. We use the Bica et al. (2020) catalogue to search for high-confidence star clusters (object class “C”) that satisfy our metallicity and age criteria. We find 14 SMC clusters fall within our definition of GC, which is slightly larger than the numbers found for NGC 6822 and IC 10.
@@
The specific frequency of GCs, defined as the number of GCs per unit of galaxy luminosity, was introduced to measure the richness of GC systems in galaxies (Harris and van den Bergh 1981) and remains commonly used today. It is calculated as , where is the absolute V-band magnitude of the host galaxy and is the number of GCs. The values are impacted by the galaxy’s recent star formation history, affecting its , and the dynamical evolution history of its GCs, affecting its . Using the galaxy absolute magnitudes from Table 1, and of 11 and 9, we find that is and in NGC 6822 and IC 10, respectively. The uncertainty in reflects both Poisson noise in the GC counts and the reported uncertainty in . These values are displayed in the top panel of Fig. 19, plotted against . This plot also shows and values for the sample of LG and isolated late-type dwarf galaxies studied by Forbes et al. (2018b), which includes some of the systems originally presented in Georgiev et al. (2010). We add further datapoints from the recent studies by Karim et al. 2024 (IC 2574) and Larsen et al. 2025 (Holmberg II), and we include values for the SMC calculated assuming both one and 14 GCs (connected via a dashed line). We see that NGC 6822 and IC 10 have high values for their luminosity, and lie at the high end of the distribution when compared to the sample of dwarfs as a whole. This is particularly true when considering that this plot only shows dwarf galaxies that have at least one identified GC; many others show none at all. Of the three late-type galaxies with higher values, one of them is the LG transition dwarf galaxy Pegasus (Cole et al. 2017) while the other two are more distant systems (Georgiev et al. 2010, KK 16 and UGC 685).
@@ 8 Conclusions
The clusters identified have masses spanning from to , and span ages from 10 Myr to 10 Gyr. IC 10 has more young clusters than NGC 6822, and its young clusters also reach higher masses (), consistent with this galaxy having a higher truncation mass on account of its higher SFR surface density. We found several examples of old massive () clusters in both dwarfs that could be the descendants of the young star clusters recently uncovered in low-mass galaxies at high redshift (e.g. Mowla et al. 2024; Adamo et al. 2024). We highlighted the existence of a particularly exceptional object in the outskirts of NGC 6822, SC7, for which we made the first mass determination of , rivalling the most massive GC in the MW. Combined with its high ellipticity and moderately high metallicity, this object bears many hallmarks of being the accreted core of a now-defunct smaller system.
@@
Lastly, we used well-defined criteria to select a subsample of clusters as candidate GCs; this yields 11 objects in NGC 6822 (folding in objects from previous studies) and 9 in IC 10. The GCs are more broadly distributed than other star clusters, with 7 of the 11 GCs lying outside the galaxy half-light radius in NGC 6822 and all 9 of the GCs in IC 10 lying outside of it. We placed the systems on the well-known GC scaling relations and found that both galaxies have high values for their luminosity, and lie at the high end of the distribution when compared to the population of dwarfs as a whole. We note that although the new star clusters uncovered in this work have increased the GC count in both galaxies, more GCs may still await detection. Indeed, the Euclid ERO images do not cover the entire halos of these systems, where more GCs may reside (including 3 previously known GCs in NGC 6822). This highlights the need for systematic, high-sensitivity, wide-field surveys of dwarf galaxies to obtain an accurate census of their GC populations.
Data Availability
Tables 2, 3, 9, 9, 11 and 11 are available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/.